# Data-Driven Goal Recognition Design for General Behavioral Agents

### v2

@@ Abstract
Goal recognition design aims to make limited modifications to decision-making environments with the goal of making it easier to infer the goals of agents acting within those environments. Although various research efforts have been made in goal recognition design, existing approaches are computationally demanding and often assume that agents are (near-)optimal in their decision-making. To address these limitations, we introduce a data-driven approach to goal recognition design that can account for agents with general behavioral models. Following existing literature, we use worst-case distinctiveness (wcd) as a measure of the difficulty in inferring the goal of an agent in a decision-making environment. Our approach begins by training a machine learning model to predict the wcd for a given environment and the agent behavior model. We then propose a gradient-based optimization framework that accommodates various constraints to optimize decision-making environments for enhanced goal recognition. Through extensive simulations, we demonstrate that our approach outperforms existing methods in reducing wcd and enhancing runtime efficiency in conventional setup. Morevoer, our approach also adapts to settings in which existing approaches do not apply, such as those involving flexible budget constraints, more complex environments, and suboptimal agent behavior. Finally, we have conducted human-subject experiments which confirm that our method can create environments that facilitate efficient goal recognition from real-world human decision-makers.
@@ 1 Introduction
With the rapid advancement of artificial intelligence (AI), there has been a surge in interest in human-AI collaboration, aiming to synergize human and AI capabilities across various domains such as gaming, e-commerce, healthcare, and workflow productivity. Designing AI agents to work alongside humans requires these agents to understand and infer human goals and intentions. While there has been abundant research in goal recognition [23, 24] that aims to infer human goals by observing their actions, this work focuses on the goal recognition design problem [12], where one needs to identify how to modify decision-making environments to enable better goal recognition.
@@
Our work is built on the goal recognition design problem formulated by Keren et al. [12]. They proposed the worst-case distinctiveness (wcd) metric, defined as the maximum number of decisions an agent can make without revealing its goal, to measure the difficulty of inferring the agent’s goal. They then aimed to modify the decision-making environment, through removing allowable actions from states, to optimize this measure. Since the introduction of this work, there have been several follow-up works that extend the formulation to deal with different settings, such as stochastic settings [28, 29] and partial observability [14]. More discussion can be found in the survey by Keren et al. [15].
@@
While there has been significant progress in goal recognition design, there are two main limitations in the literature. First, current approaches require evaluating the difficulty of goal recognition, i.e., worst-case distinctiveness (wcd), for a large number of potential modifications to the decision-making environment. Since each evaluation of wcd requires solving the optimal policy multiple times to each of the goals, this process is computationally demanding and limits the scalability of these approaches, especially in virtual domains where the number of possible environment configurations is large. Second, decision-making agents in goal recognition design are often assumed to be optimal. While there are efforts addressing suboptimal agents [13, 27], these primarily focus on settings where there is a limited number of deviations from optimal decision-making. This assumption of optimality is limiting in scenarios involving human agents, who are known to often systematically deviate from optimal decision-making due to cognitive and informational constraints [10, 19].
@@
To address these limitations, we propose a framework for data-driven goal recognition design with general agent behavior. To relax the optimality assumption, we explicitly incorporate models of agent behavior into the optimization framework to better represent behavioral agents. To tackle the computational challenges, our approach leverages data-driven methods for goal recognition design. The core idea involves building a machine learning oracle that predicts the difficulty of goal recognition (e.g., wcd) given a decision-making environment and an agent’s behavioral model. This oracle is trained on datasets generated from simulations, allowing for the evaluation of wcd in any given environment and agent model. Such an approach significantly accelerates the evaluation of wcd during run-time optimization. Once the machine learning oracle is established, we apply the general gradient-based optimization approach to minimize wcd, employing Lagrangian relaxation to manage constraints on environment modifications. This allows us to address more general forms of objectives and constraints that existing approaches in the literature cannot address.
@@
To evaluate our framework, we conducted a series of simulations and human-subject experiments. We start with simulations in the standard setup, with the grid world environment and optimal agent assumption. We show that our approach outperforms existing baselines in reducing worst-case distinctiveness (wcd) and demonstrates considerably better run-time efficiency. We then conducted additional simulations to showcase that our approach can generalize to settings beyond the capabilities of existing approaches in the literature. These include scenarios involving flexible budget constraints, more complex environments, and suboptimal agent behavior. Lastly, we have conducted human-subject studies demonstrating that our method can be leveraged to design environments that facilitate efficient goal recognition from real-world human decision-makers. The results highlight the potential of our approach to enable more efficient human-AI collaboration.
@@
    We propose a data-driven optimization framework for goal recognition design. This framework comprises a predictive model that estimates the wcd for a given environment and a model of agent behavior. We then utilize a gradient-based optimization method for goal recognition design. The framework is flexible to accommodate various environments, design spaces, and agent behavior.
@@
    Through extensive simulations, we show that our framework not only outperforms existing approaches in goal recognition design within standard settings but also adapts to scenarios that existing approaches cannot handle, including general optimization criteria, complex environments, and suboptimal agent behavior.
@@
    Through human-subject experiments, we demonstrate that our approach can adapt to agent models trained on human behavior. Furthermore, our framework can create environments that enable more effective goal recognition with real-world humans. To the best of our knowledge, our work is the first to utilize human-subject experiments to evaluate the environment design for goal recognition.
@@ 2 Related Work
Our work contributes to the expanding field of human-AI collaboration. Recent research has indicated that optimizing AI alone is insufficient for maximizing the performance of human-AI teams [1, 2]. To develop truly collaborative AI agents, it is crucial to equip them with the ability to comprehend and predict the intentions and goals of their human counterparts. This challenge lies at the heart of goal recognition research [11, 23, 9, 24, 20, 17]. Our work specifically focuses on goal recognition design, an extension of goal recognition that includes modifying environments to better facilitate the process of recognizing goals.
@@
Goal recognition design was formulated by Keren et al. [12]. Since this seminal work, numerous research efforts have expanded the concept to accommodate stochastic environments [29, 27], different levels of observability [14, 30], and a variety of design spaces [18]. The studies most closely aligned with our approach focus on suboptimal agents [13, 27]. However, these studies characterize suboptimality by limiting deviations from the optimal policy, a method that may not adequately represent behavior of human agents, who frequently deviate systematically from optimal decision-making. Additionally, most existing work requires evaluating the difficulty of goal recognition in run time for numerous environmental modifications, which constrains the scalability of these methods. Our work sets itself apart by broadening the goal recognition design question to incorporate general agent behavioral models and by implementing a data-driven optimization approach.
From a technical standpoint, we adopt a data-driven optimization approach for goal recognition design. The use of data-driven tools in optimization has gained increasing prominence in the field of mechanism design, as evidenced by a range of recent studies [5, 7, 8, 4, 16, 22, 21, 3, 32]. Moreover, our work closely aligns with the recent studies that explicitly encoded human behavior and responses in the design of computational and learning systems [25, 26, 31, 32, 6]. While our work shares a similar motivation, our approach and problem formulation differ from theirs.
@@ 3.1 Problem Setting
We formulate the decision-making environment and behavioral agent models. We also explain the commonly-used metrics for evaluating goal recognition difficulty.
@@
Models of behavioral agents. We represent the agent’s decision-making policy in a general form $\Pi:S\times T\rightarrow A$. Specifically, for an agent with a decision-making policy $\pi\in\Pi$, the agent will execute the action $\pi(s,t)$ in state $s$ at time $t$. The agent is conceptualized as a planner $H:W\rightarrow\Pi$, where the input is an environment $w\in W$, and the output is a policy $\pi\in\Pi$. To illustrate our formulation, consider a agent parameterized by a time-variant discounting function $d(t)$. The standard optimal agent model corresponds to a fixed discounting factor $\gamma\in(0,1]$ with $d(t)=\gamma^{t}$. We can also represent an agent with present bias [19] by adopting a hyperbolic discounting factor $d(t)=\frac{1}{1+kt}$, where $k>0$. We would like to highlight that our approach not only accounts for standard analytical closed-form expression of agent behavior. It can also accounts for scenarios where an agent policy $\pi$ is a machine learning model, i.e., a neural network trained on human behavioral data.
Worst-case distinctiveness (wcd). To evaluate the difficulty of goal recognition, we adhere to standard literature and focus on the measure of worst-case distinctiveness (wcd[12], defined as the maximum number of steps an agent can take before revealing its goal. To compute the wcd for a given suboptimal agent $h\in H$ in an environment $w\in W$, we evaluate the path for the agent to each goal and compute the number of actions from the initial state that are identical for every goal.
3.2 Goal Recognition Design with General Behavioral Agents
We formalize the goal recognition design problem with general behavioral agents. Given an environment $w$ and an agent with a behavior model $h$, we denote the worst-case distinctiveness of environment $w$ for agent $h$ as $\textit{wcd}(w,h)$. Each type of modification $1\leq i\leq N$ will incur a cost $c_{i}(w,w^{\prime})$ that must fall in budget constraint $B_{i}$. The objective of the goal recognition design problem is to alter the environment from $w$ to $w^{\prime}$ in a way that minimizes $\textit{wcd}(w^{\prime},h)$, while satisfying the constraint that the cost of the modifications does not exceed the budget.
@@ 3.3 Our Proposed Method
Existing approaches in goal recognition design require evaluating wcd for a large number of potential modifications to the environments to identify the optimal modifications. Since evaluating wcd requires evaluating the agent policy for multiple goals, it presents significant computational overhead and limits the scalability of these approaches. To overcome this challenge, we propose leveraging machine learning to expedite run-time computations. The main idea, as summarized in Figure 1, is to first train a machine learning model that predicts wcd for any given pair of decision-making environment $w$ and agent behavioral model $h$. After obtaining this machine learning model and utilizing its differentiable property, we develop an optimization framework that applies gradient-based optimization methods to the Lagrangian relaxation of the constrained optimization formulation in (1).
The main benefits of our approach compared with existing methods include: 1) the ability to incorporate different agent models $h$, 2) flexibility for various forms of optimization objectives and constraints, and 3) run-time efficiency.
@@
Figure 1: We first train a predictive model to predict wcd from simulated data. We then perform gradient-based optimization that leverages the predictive model to identify environment modifications that minimize wcd with a given model of agent behavior.
@@
Figure 2: The example showcases two benchmark environments. The first, displayed on the left, is a grid world environment. The agent starts at a position marked ’S’ and aims to reach one of the goal positions labeled ’G’. The agent must navigate through the grid, avoiding blocked cells marked with ’x’. The second environment, illustrated on the right, is an Overcooked-AI setting. In this scenario, the agent’s objective is to pick up ingredients and complete the cooking of their target recipe, which constitutes their goal.
Predictive model for wcd. To build the predictive model for wcd, we curate a training dataset through simulations. For a given environment $w$ and agent behavioral model $h$, we can obtain $\textit{wcd}(w,h)$ by solving for the agent’s actions towards different goals. After collecting a training dataset, we train the predictive model using a convolutional neural network. Implementation details are included in the appendix.
Optimization procedure. After obtaining the predictive model, we develop a gradient-based optimization framework. This framework strictly generalizes the existing literature in goal recognition design where the space of modifications are limited (e.g., usually limits to blocking an action in MDP) [15].
The first step involves transforming the constrained optimization problem in (1) into an unconstrained optimization problem using Lagrangian relaxation:
We then perform gradient descent on the relaxed Lagrangian. As environment modifications are often discrete (e.g., whether to block a cell in the grid world), we apply a discrete gradient descent procedure. Specifically, at each step of gradient descent, we obtain a gradient, which is a vector indicating the suggested change magnitude for each element (such as a cell in the grid world). We then select the element with the highest gradient value and make the corresponding change. Note that some suggested modifications may not be valid; for instance, we cannot block a cell that is already blocked in the grid world. In such cases, we proceed to the element with the next highest gradient value, continuing this process until a valid modification is made. This modification procedure is repeated until the gradient descent converges. Note that with this Lagrangian relaxation, we cannot directly set the budget, however, based on duality, a larger Lagrangian multiplier $\lambda$ corresponds to a smaller budget $B$, in the original constrained optimization formulation. In our experiments, instead of selecting different budgets directly, we choose varying Lagrangian multipliers and record the realized costs of the modifications.
4 Experiment Setup
4.1 Benchmark Domains
We utilize two benchmark domains. The first is the standard basic grid world environment, commonly used in goal recognition design literature. The second is the Overcooked-AI environment [2], an complex environment with a richer set of environment modification. This environment is particularly relevant to the downstream implications of our work, namely in supporting human-AI collaboration.

### v3

# Goal Recognition Design for General Behavioral Agents using Machine Learning

@@ Abstract
Goal recognition design (GRD) aims to make limited modifications to decision-making environments to make it easier to infer the goals of agents acting within those environments. Although various research efforts have been made in goal recognition design, existing approaches are computationally demanding and often assume that agents are (near-)optimal in their decision-making. To address these limitations, we leverage machine learning methods for goal recognition design that can both improve run-time efficiency and account for agents with general behavioral models. Following existing literature, we use worst-case distinctiveness (wcd) as a measure of the difficulty in inferring the goal of an agent in a decision-making environment. Our approach begins by training a machine learning model to predict the wcd for a given environment and the agent behavior model. We then propose a gradient-based optimization framework that accommodates various constraints to optimize decision-making environments for enhanced goal recognition. Through extensive simulations, we demonstrate that our approach outperforms existing methods in reducing wcd and enhances runtime efficiency. Moreover, our approach also adapts to settings in which existing approaches do not apply, such as those involving flexible budget constraints, more complex environments, and suboptimal agent behavior. Finally, we conducted human-subject experiments that demonstrate that our method creates environments that facilitate efficient goal recognition from human decision-makers.
@@ 1 Introduction
With the rapid advancement of artificial intelligence (AI), there has been a surge in interest in human-AI collaboration, aiming to synergize human and AI capabilities across various domains, such as gaming, e-commerce, healthcare, and workflow productivity. Designing AI agents to work alongside humans requires these agents to understand and infer human goals and intentions. While there has been abundant research in goal recognition (Ramírez & Geffner, 2010; Sukthankar et al., 2014), aiming to infer human goals by observing their actions, this work focuses on the goal recognition design problem (Keren et al., 2014), which seeks to identify how to modify decision-making environments to make it easier to infer human goals.
@@
Our work is built on the GRD problem formulated by Keren et al. (2014). They proposed the worst-case distinctiveness (wcd) metric, defined as the maximum number of decisions an agent can make without revealing its goal, to measure the difficulty of inferring the agent’s goal. They then aimed to modify the decision-making environment, by removing allowable actions from states, to optimize this measure. Since the introduction of this work, there have been several follow-up works to deal with different settings, such as stochastic settings (Wayllace et al., 2016; 2017) and partial observability (Keren et al., 2016). More discussion can be found in the survey by Keren et al. (2021).
@@
While there has been significant progress in GRD, there are two main limitations in the literature. First, existing approaches often assume that decision-making agents are optimal or near-optimal decision makers (Keren et al., 2015; Wayllace & Yeoh, 2022). This assumption is unrealistic with human decision makers, who are known to often systematically deviate from optimal decision-making due to cognitive and informational constraints (Kahneman, 2003; O’Donoghue & Rabin, 1999). Second, existing approaches require evaluating the difficulty of goal recognition, i.e., worst-case distinctiveness (wcd), for a large number of modifications to the environment. This computational requirement limits the scalability of these approaches, especially for virtual environments when there might be frequent updates of the environment, e.g., e-commerce platforms might want to frequently adjust the website design to enhance user intent inference.
@@
We propose a framework for GRD with general agent behavior to address these limitations. To relax the optimality assumption, we explicitly incorporate models of agent behavior into the optimization framework. To tackle the computational challenges, our approach leverages data-driven methods for goal recognition design. The core idea is to build a machine learning oracle that predicts the difficulty of goal recognition (e.g., wcd) given a decision-making environment and an agent’s behavioral model. This oracle is trained on datasets generated from simulations, enabling efficient evaluation of wcd for any given environment and agent model. Such an approach significantly accelerates the evaluation process during optimization. Once the machine learning oracle is established, we apply a general gradient-based optimization method to minimize wcd, using Lagrangian relaxation to handle constraints on environment modifications. Beyond addressing the two main limitations, this optimization procedure also supports more general forms of objectives and constraints that existing approaches in the literature cannot accommodate.
@@
To evaluate our framework, we conducted a series of simulations and human-subject experiments. We start with simulations in the standard setup in the literature, with the grid world environment and optimal agent assumption. We show that our approach outperforms existing baselines in reducing worst-case distinctiveness (wcd) and achieves considerably better run-time efficiency. We then conducted additional simulations to demonstrate that our approach can generalize to settings that existing methods cannot address, including scenarios involving flexible budget constraints, more complex environments, and suboptimal agent behavior. Lastly, we have conducted human-subject studies demonstrating that our method can be leveraged to design environments that facilitate efficient goal recognition from real-world human decision-makers. The results highlight the potential of our approach to enable more efficient human-AI collaboration.
@@
    We propose a data-driven optimization framework that accommodates general agent behavior and improves the efficiency of goal recognition design, addressing two main limitations in the existing literature while enabling a more flexible optimization procedure. The framework consists of a predictive model that estimates the wcd for a given environment and a model of agent behavior. We then apply a gradient-based optimization method to perform goal recognition design. The framework is runtime efficient, capable of incorporating general agent behavior, and adaptable to different design spaces and constraints.
@@
    Through extensive simulations, we show that our framework not only outperforms existing approaches in goal recognition design, both in reducing worst-case distinctiveness wcd and improving run-time efficiency in standard settings, but also adapts to scenarios that existing methods cannot handle, including general optimization criteria, complex environments, and suboptimal agent behavior.
@@
    Through human-subject experiments, we demonstrate that our approach adapts to settings in which decision-making agents are real-world humans. To the best of our knowledge, our work is the first to evaluate the environment design for goal recognition with human-subject experiments.
@@ 2 Related Work
Our work contributes to the field of human–AI collaboration. Recent research shows that optimizing AI alone is insufficient to maximize the performance of human–AI teams (Bansal et al., 2021; Carroll et al., 2019; Yu et al., 2024), motivating efforts to understand and model human behavior in interactions with AI systems (Choudhury et al., 2019; Shah et al., 2019; Kwon et al., 2020; Chan et al., 2021; Reddy et al., 2021; Narayanan et al., 2022a; b; 2023; Treiman et al., 2023; 2024; 2025). To build truly collaborative AI agents, they must also anticipate the intentions and goals of their human counterparts. This challenge lies at the core of goal recognition research (Kautz, 1991; Ramírez & Geffner, 2010; Hong, 2001; Sukthankar et al., 2014; Pereira et al., 2017; Masters et al., 2021). Our work focuses on goal recognition design, an extension of goal recognition that incorporates environment modifications to make goals easier to recognize.
@@
Goal recognition design (GRD) was formulated by Keren et al. (2014). Since this seminal work, numerous research efforts have extended to accommodate stochastic environments (Wayllace et al., 2017; Wayllace & Yeoh, 2022), varying levels of observability (Keren et al., 2016; Wayllace et al., 2020), and diverse design spaces (Mirsky et al., 2019). The studies most closely aligned with our approach focus on suboptimal agents (Keren et al., 2015; Wayllace & Yeoh, 2022). However, these studies characterize suboptimality by limiting deviations from the optimal policy, a method that may not adequately capture the behavior of human agents, who frequently deviate systematically from optimal decision-making. Moreover, most existing work requires evaluating goal recognition difficulty at run time across a large number of environmental modifications, which significantly limits scalability. Our work distinguishes itself by broadening GRD to incorporate general models of agent behavior and by implementing a data-driven optimization approach.
There has been effort to accelerate GRD by using heuristic approaches that avoid the prohibitively large search space of possible modifications. In their seminal work, Keren et al. (2014) proposed the pruned-reduce method, which restricts modifications to action removal and leverages a key insight: blocking actions do not introduce new goal-directed paths. This alleviates the computational burden of exhaustive search. More recent work by Pozanco et al. (2024) accelerates GRD by using a plan library from a top-quality planner (Katz et al., 2020) to guide initial heuristic exploration. Similarly, Au (2024) expanded the design space with block-level modifications—simultaneous, interconnected changes—to improve efficiency. While these approaches make notable strides, they still rely on exhaustive search and repeated computation of complex metrics, which remains the primary computational bottleneck. In contrast, our method adopts a data-driven optimization framework. Rather than evaluating the metric (wcd) at each step of the optimization process, we train predictive models to approximate it directly, enabling rapid evaluation and comparison of design candidates. Moreover, the differentiability of the predictive model allows for gradient-based optimization in place of exhaustive search. Together, these advances significantly reduce computational overhead and open new possibilities for scaling GRD to more complex settings.
Our methodology aligns with recent advances in data-driven optimization for mechanism design (Dütting et al., 2019; Golowich et al., 2018; Curry et al., 2020; Kuo et al., 2020; Rahme et al., 2020; Peri et al., 2021; Cornelisse et al., 2022; Yu et al., 2023), which leverage machine learning to accelerate otherwise intractable tasks. It also aligns with recent research that incorporates human models into learning and optimization processes (Ho et al., 2016; Evans et al., 2016; Shah et al., 2019; Tang & Ho, 2019; 2021; Carroll et al., 2019; Tang et al., 2021a; b; Yu & Ho, 2022; Feng et al., 2024; Kasumba et al., 2025). While we share the high-level motivation, our problem formulation and methodological choices offer novel contributions and advance the study of goal recognition design through a principled integration of predictive modeling and optimization.
@@ 3.1 Problem Setting
Models of behavioral agents. We represent the agent’s decision-making policy in a general form $\Pi:S\times T\rightarrow A$. Specifically, for an agent with a decision-making policy $\pi\in\Pi$, the agent will execute the action $\pi(s,t)$ in state $s$ at time $t$. The agent is conceptualized as a planner $H:W\rightarrow\Pi$, where the input is an environment $w\in W$, and the output is a policy $\pi\in\Pi$. To illustrate our formulation, consider an agent parameterized by a time-variant discounting function $d(t)$. The standard optimal agent model corresponds to a fixed discounting factor $\gamma\in(0,1]$ with $d(t)=\gamma^{t}$. We can also represent an agent with present bias (O’Donoghue & Rabin, 1999) by adopting a hyperbolic discounting factor $d(t)=\frac{1}{1+kt}$, where $k>0$. Note that our approach not only accounts for standard analytical closed-form expression of agent behavior. It can also account for scenarios where an agent policy $\pi$ is a machine learning model, i.e., a neural network trained on human behavioral data.
Worst-case distinctiveness (wcd). To evaluate the difficulty of goal recognition, we follow the standard literature and focus on worst-case distinctiveness (wcd(Keren et al., 2014), defined as the maximum number of steps an agent can take that are consistent with multiple goals before its true goal becomes distinguishable. In other words, it measures the number of initial steps that overlap across trajectories toward different goals. To compute the wcd for a given agent $h\in H$ in an environment $w\in W$, we evaluate the path for the agent to each goal and compute the number of actions from the initial state that are identical for each goal. While we focus on wcd in this work, our method is agnostic to the specific metric used, as long as the metric is computable, a predictive model can be trained to approximate it and enable efficient evaluation.
3.2 Goal Recognition Design (GRD) Formulation
We formalize the GRD problem with general behavioral agents. Given an environment $w$ and an agent with a behavior model $h$, we denote the worst-case distinctiveness of environment $w$ for agent $h$ as $\textit{wcd}(w,h)$. Each type of modification $1\leq i\leq N$ will incur a cost $c_{i}(w,w^{\prime})$ that must fall in budget constraint $B_{i}$. The objective of the GRD problem is to alter the environment from $w$ to $w^{\prime}$ in a way that minimizes $\textit{wcd}(w^{\prime},h)$, while satisfying the constraint that the cost of the modifications does not exceed the budget.
@@ 3.3 Our Proposed Method
Existing approaches to GRD require evaluating wcd for a large number of potential environment modifications in order to identify the optimal ones. Since computing wcd involves evaluating the agent’s policy across multiple goals, it introduces significant computational overhead and relies on strong assumptions about agent behavior. To address this challenge, we propose leveraging machine learning to expedite runtime computation and explicitly account for general agent models. The main idea, illustrated in Figure 1, is to first train a machine learning model that predicts wcd for any given pair of decision-making environment $w$ and agent behavioral model $h$. Once this predictive model is trained, we exploit its differentiability to develop an optimization framework that applies gradient-based methods to the Lagrangian relaxation of the constrained optimization problem defined in Equation 1. This approach not only addresses the two key limitations—computational challenges and restrictive assumptions about agent behavior—but also provides flexibility for accommodating various forms of optimization objectives and constraints.
@@
Figure 1: We first train a wcd predictor from simulated data. We then perform gradient-based optimization that leverages the predictor to identify environment modifications that minimize wcd with a given agent behavior model.
Predictive model for wcd. To build the predictive model for wcd, we curate a training dataset through simulations. For an environment $w$ and agent behavioral model $h$, we can obtain $\textit{wcd}(w,h)$ by solving for the agent’s actions towards different goals. After collecting a training dataset, we train the predictive model using a convolutional neural network. The implementation details are in Section 5.1.1 and the appendix.
Optimization procedure. After obtaining the predictive model, we develop a gradient-based optimization framework. This framework generalizes the existing literature in GRD where the space of modifications is limited (e.g., usually limits to blocking an action in MDP) (Keren et al., 2021). The first step of the optimization procedure involves transforming the constrained optimization problem in equation 1 into an unconstrained optimization problem using Lagrangian relaxation:
We then perform gradient descent on the relaxed Lagrangian with respect to $w$. As environment modifications are often discrete (e.g., to block a cell in the grid world), we apply a discrete gradient descent procedure. Specifically, at each descent step, we obtain a gradient, which is a vector indicating the suggested change magnitude for each element (such as a cell in the grid world). We then select the element with the gradient value of the largest magnitude and make the corresponding change. Note that some suggested modifications may be invalid; for instance, we cannot block an already blocked cell in the grid world. In such cases, we proceed to the element with the next highest gradient value, continuing this process until a valid modification is made. This modification procedure is repeated until the gradient descent converges. Note that with this Lagrangian relaxation, while we cannot directly set the budget, based on duality, a larger Lagrangian multiplier $\lambda$ corresponds to a smaller budget $B$ in the original constrained optimization formulation. In practice, designer can choose to vary the size of the Lagrangian multiplier to modulate the budget. In our experiments, we choose varying Lagrangian multipliers and record the realized costs of the modifications.
4 Experiment Setup
4.1 Benchmark Domains
We utilize two benchmark domains. The first is the standard basic grid world environment, commonly used in the GRD literature. The second is the Overcooked-AI environment (Carroll et al., 2019), a more complex environment with a richer set of environment modifications. This environment is particularly relevant to the downstream implications of our work, namely in supporting human-AI collaboration.
@@ 4.1 Benchmark Domains
Figure 2: The benchmark environments. The left one is a grid world. The agent starts at a position marked ’S’ and aims to reach one of the goal positions labeled ’G’. The agent must navigate through the grid, avoiding blocked cells marked with ’x’. The right one is an Overcooked-AI setting, where the agent’s objective is to pick up ingredients and complete their target recipe, which constitutes their goal.